Find the area of the shaded region. 6- (0,4) y=4 - x2 LY X 2 (3,-5) O A. 3 ОВ. 5 3 O C. 5 OD 23 3
Answer #1

y g. 4-772 Y = 4-02 We have to find Area of shaded region. 0 (3-5) 2 4-22 Area = M dy dx + 1 l dyda Oy=0 y = 4 22 + ( 28 47 3